The summaries of data from the balance sheet, income statement, and retained earnings statement for two corporations, Walco Corporation and Gunther Enterprises, are presented below for 2017 Determine the missing amounts. Assume all changes in stockholders equity are due to changes in retained earnings Walco Corporation $117,000 72,000 Gunther Enterprises Beginning of year Total assets $159,000 Total liabilities Total stockholders equity 71,300 End of year 182,000 Total assets Total liabilities Total stockholders equity 120,000 59,000 51,900 Changes during year in retained earnings Dividends 4.600 Total revenues 215,000 Total expenses 169,000 80,000
Answer #1
Walco Corp Gurther Ent A Total Assets=Total Liabilities+Total Sharholders Equity
Beginning of Year 117000=72000+A
Total Assets        1,17,000        1,59,000 A=45000
Total Liabilities           72,000           87,700
D Total Assets=Total Liabilities+Total Sharholders Equity
Total Stockholder's Equity           45,000           71,300 159000=D+71300
D=87700
End of Year
Total Assets        1,71,900        1,82,000 B Total Assets=Total Liabilities+Total Sharholders Equity
Total Liabilities        1,20,000           59,000 B=120000+51900
B=171900
Total Stockholder's Equity           51,900        1,23,000
E Total Assets=Total Liabilities+Total Sharholders Equity
Changes during year in REs: 182000=59000+E
Dividends           39,100              4,600 E=123000
Total Revenues        2,15,000        1,36,300 C Total Revenues 2,15,000
Total Expenses        1,69,000           80,000 Total Expense 1,69,000
Income During the Year       46,000
Change in Stockholder's Equity-Increase         6,900 (51900-45000)
Dividend Paid       39,100
F Total Expense       80,000
Dividend Paid         4,600
Change in Stockholder's Equity-Increase       51,700 (123000-71300)
Total Revenue 1,36,300